Solve using the quadratic formula. 5x2 + 13x = 6
Answer #1

# 572 +132 = 6 o § 5712 + 133-6 = apply - 13+ J(13)2 - 4x 5X(-6) Soridharacharya -6+ Jb2-490 sule a equan?+ba+c=0 2a : x = 2X

If any doubt please comment